OpenStudy (anonymous):

Using the formula P = 2(a + b), find a when P = 15 and b = 5.

OpenStudy (anonymous):

substitute instead of each letter (variable) its number so its gonna be 15=2(a+5)--> can u solve it now ?

OpenStudy (anonymous):

15 = 2(a+5) 15= 2a+10 15 - 10 5 = 2a 5/2 = 2.5 a = 2.5

OpenStudy (anonymous):

Sorry Eyad! Didn't look before posting :(

OpenStudy (anonymous):

nvm @musicalrose : u did great :)